OpenStudy (anonymous):

If you rent a car for one day and drive it for 100 miles, the cost is $40.00. If you drive it 220 miles the cost is $46. Use the linear function to find out how much it will cost to rent the car for one day if you drive it 300 miles.

OpenStudy (anonymous):

y = 1/20x + 35 Now, if x = 300 miles, then y= 1/20*300 + 35 y = 15 + 35 y = 50 It is like each mile costs 5 cents, and the price to rent the car was $35. Got it?